assume that a procedure yields a binomial distribution with a trial repeated n times. use the binomial probability formula to find the probability of x successes given the probability p of success on a single trial

n=16, x=12,p=0.75

p(12)=

How do I use my calculator? What is the ! on the calculator?  can u use stat cruch for this problem?
Answer #1

Solution :

Given that ,

p = 0.75

1 - p = 1 -0.5 = 0.25  

n = 16

x = 12

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * (1 - p)n - x

P(X = 12 ) = (16! / (16 - 12)!) * 0.7512 * 0.25)16 - 12

= 0.2252

Probability = 0.2252